OpenStudy (anonymous):

tn=8.5n-30 Show me step by step??

OpenStudy (anonymous):

whats the 7th term in this sequence?

hero (hero):

t(n) = 8.5n - 30 t(7) = 8.5(7) - 30

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So its basically y=mx+b?

hero (hero):

yup

OpenStudy (anonymous):

Thanks :)